# On the Non-Identifiability of Steering Vectors in Large Language Models

### 05 Mar 2026 v3

@@ C.1 Geometric Intuition for Null-Space Ambiguity
Consider a simplified example where a 3D steering vector $v\in\mathbb{R}^{3}$ affects 2D outputs $o\in\mathbb{R}^{2}$ through a projection matrix $J\in\mathbb{R}^{2\times 3}$. By the rank-nullity theorem, the null space $\ker(J)$ is a 1D subspace since $\dim(\ker(J))=3-\text{rank}(J)=3-2=1$.
@@
Figure 5: Geometric intuition for null-space ambiguity. The output depends only on the $(v_{1},v_{2})$ projection (row space of $J$), while $v_{3}\in\ker(J)$ is invisible; hence $J(v+\alpha v_{3})=Jv$.
For concreteness, suppose $J=[I_{2}\mid 0]$ projects onto the first two coordinates. Then $\ker(J)=\text{span}\{(0,0,1)\}$ is the $v_{3}$ axis. The key insight generalizes: for any projection matrix $J$, directions in $\ker(J)$ are invisible to outputs. Figure 5 illustrates this geometric intuition.
Any vector of the form $v^{\prime}=v+\alpha v_{3}$ for $\alpha\in\mathbb{R}$ produces identical outputs:
$J(v+\alpha v_{3})$ $=Jv+\alpha Jv_{3}$ (30)
$=Jv+\alpha\cdot 0\quad(v_{3}\in\ker(J))$
$=Jv.$
Since $\alpha$ can take infinitely many values and $v^{\prime}\not\propto v$ for generic choices of $\alpha$, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
C.2 Why More Data Doesn’t Help
Common misconception: "If we collect more steering examples, we can pin down the unique vector."
Why this fails:
Consider observing steering effects on $N$ prompts $\{x_{1},\ldots,x_{N}\}$. More fundamentally, any $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v_{0}=0$ by definition since this is a property of the model weights and not the prompt distribution. Stacking observations across prompts therefore cannot remove directions from $\ker(J_{\ell})$. Each observation provides:
$o_{i}=J_{\ell}(x_{i})v+\eta_{i}$ (31)
Stacking these:
$\begin{bmatrix}o_{1}\\ o_{2}\\ \vdots\\ o_{N}\end{bmatrix}=\begin{bmatrix}J_{\ell}(x_{1})\\ J_{\ell}(x_{2})\\ \vdots\\ J_{\ell}(x_{N})\end{bmatrix}v+\begin{bmatrix}\eta_{1}\\ \eta_{2}\\ \vdots\\ \eta_{N}\end{bmatrix}$ (32)
The stacked Jacobian $J_{\text{stack}}\in\mathbb{R}^{(N\cdot V)\times d}$ has null space:
$\ker(J_{\text{stack}})=\bigcap_{i=1}^{N}\ker(J_{\ell}(x_{i}))$ (33)
@@ C.2 Why More Data Doesn’t Help
Figure 6: Illustration of why prompt diversity does not resolve null-space ambiguity. Each prompt induces a Jacobian that measures the same effective subspace, leaving the null-space intersection unchanged.
Critical observation: If all $J_{\ell}(x_{i})$ share a common null space (e.g., all prompts probe similar aspects of the model), then:
$\ker(J_{\text{stack}})=\ker(J_{\ell}(x_{1}))\neq\{0\}$ (34)
This means adding more prompts does not reduce the null space when all prompts probe the same effective subspace—the ambiguity persists. Figure 6 illustrates this phenomenon.
C.3 Why Orthogonal Perturbations Preserve Semantics
Setup: We test steering with $v$ versus $v+v_{\perp}$, where $v_{\perp}\perp v$ and $\|v_{\perp}\|=1$. Empirically, adding a random orthogonal direction produces nearly equivalent semantic effects. Why?
Decomposing the perturbation
Any perturbation $v_{\perp}$ can be decomposed into observable and invisible components:
$v_{\perp}=v_{\perp,\text{row}}+v_{\perp,\text{null}}$ (35)
where $v_{\perp,\text{row}}\in\operatorname{row}(J)$ (observable through outputs) and $v_{\perp,\text{null}}\in\ker(J)$ (invisible to outputs).
For a random $v_{\perp}\perp v$, the expected fraction in the null space is:
$\mathbb{E}[\|v_{\perp,\text{null}}\|^{2}]\approx\frac{\dim(\ker(J))}{d}\approx 0.20{-}0.25$ (36)
This means $\sim$20–25% of the perturbation is automatically invisible to model outputs.
Observable effect is diffuse and unstructured
The actual output change from adding $v_{\perp}$ is:
$J(v+v_{\perp})=Jv+Jv_{\perp,\text{row}}$ (37)
While $Jv_{\perp,\text{row}}$ is not necessarily small in norm, it represents a random direction in the $\sim$3200-dimensional row space of $J$. In contrast, $Jv$ is a structured, semantically coherent steering effect (e.g., consistently shifting outputs toward “honesty” or “sycophancy”).
The key distinction is not magnitude but semantic coherence: $Jv$ produces aligned, directional changes across the vocabulary, while $Jv_{\perp,\text{row}}$ produces diffuse, incoherent perturbations that average out across tokens and do not systematically shift semantic meaning in any consistent direction.
Analogy: If $Jv$ is a strong wind blowing consistently north, then $Jv_{\perp,\text{row}}$ is turbulence—it may have comparable energy but lacks directional coherence, so the overall trajectory remains northward.
Appendix D Prompt Construction Details
We construct contrastive prompt pairs for each semantic trait using template-based generation. Table 4 provides concrete examples for all three traits used in the empirical validation.
Table 4: Example contrastive prompt pairs for each semantic trait.
@@ Appendix D Prompt Construction Details
Humor Tell a lighthearted story about an embarrassing mistake. Provide a factual account of an embarrassing mistake.
@@
Appendix E Mathematical Derivations
E.1 Dimension of Observational Equivalence Class
Setup: For steering vector $v\in\mathbb{R}^{d}$ and Jacobian $J\in\mathbb{R}^{V\times d}$ with rank $r$, consider:
$v]=\{v^{\prime}\in\mathbb{R}^{d}:Jv^{\prime}=Jv\}$ (38)
Question: How many degrees of freedom exist in the equivalence class of observationally equivalent steering vectors?
Analysis: The equivalence class is an affine subspace:
$v]=v+\ker(J)$ (39)
where $\ker(J)$ is a (dr)(d-r)-dimensional linear subspace.
Parameterization: Let $\{u_{1},\ldots,u_{d-r}\}$ be an orthonormal basis for $\ker(J)$. Then:
$v]=\left\{v+\sum_{i=1}^{d-r}\alpha_{i}u_{i}:\alpha_{i}\in\mathbb{R}\right\}$ (40)
The equivalence class has (dr)(d-r) degrees of freedom.
Scaling ambiguity: Additionally, $v$ and $cv$ produce equivalent outputs (up to rescaling $\alpha$). The equivalence class modulo scaling is a (dr)(d-r)-dimensional projective space.
Illustrative example. Suppose a model has hidden size $d=4096$ and an effective Jacobian rank $r\approx 3100$. Then the equivalence class dimension is $d-r\approx 996$.
Interpretation: For every identifiable parameter (row-space direction), there are $996/3100\approx 0.32$ unidentifiable parameters (null-space directions). The under-determination ratio is approximately $1:3$.
E.2 Fisher Information and Cramér-Rao Bound
Setup: Consider the statistical model:
$o(x;v,\alpha)\sim p(o|x,v,\alpha)$ (41)
where $v$ is the parameter to estimate.
Question: Can we achieve better identifiability by collecting more samples?
Fisher Information Matrix: For the linear Gaussian model $o=Jv+\eta$ with $\eta\sim\mathcal{N}(0,\sigma^{2}I)$:
$\mathcal{I}(v)$ $=\mathbb{E}_{o}\Bigg[\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)$ (42)
$\qquad\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)^{\!\top}\Bigg]$
$=\frac{1}{\sigma^{2}}J^{\top}J.$
Cramér-Rao Lower Bound: The covariance of any unbiased estimator $\hat{v}$ satisfies:
$\text{Cov}(\hat{v})\succeq\mathcal{I}(v)^{-1}=\sigma^{2}(J^{\top}J)^{+}$ (43)
where $(J^{\top}J)^{+}$ is the Moore-Penrose pseudo-inverse.
Implications: For null-space directions $u_{i}\in\ker(J)$, the Fisher information is degenerate:
$u_{i}^{\top}(J^{\top}J)u_{i}=0$ (44)
Therefore, the variance of any unbiased estimator along null-space directions is unbounded:
$\text{Var}(u_{i}^{\top}\hat{v})\geq u_{i}^{\top}\mathcal{I}(v)^{-1}u_{i}=\infty$ (45)
Conclusion: The Cramér-Rao bound is infinite for null-space components, confirming that no finite amount of data can resolve the ambiguity. Non-identifiability is fundamental, not a small-sample problem. More data cannot help because the information geometry has infinite uncertainty in null-space directions.

### 01 Apr 2026 v4

@@ F.1 Geometric intuition for null-space ambiguity
Consider a simplified setting where a 3D steering vector $v\in\mathbb{R}^{3}$ affects 2D outputs $o\in\mathbb{R}^{2}$ through a projection matrix $J\in\mathbb{R}^{2\times 3}$. By the rank-nullity theorem, $\dim(\ker(J))=3-2=1$. Concretely, if $J=[I_{2}\mid 0]$ projects onto the first two coordinates, then $\ker(J)=\operatorname{span}\{(0,0,1)^{\top}\}$. Any vector of the form $v^{\prime}=v+\beta(0,0,1)^{\top}$ produces identical outputs since:
$J(v+\beta e_{3})=Jv+\beta Je_{3}=Jv+\beta\cdot 0=Jv$ (35)
The output depends only on the projection of $v$ onto $\operatorname{row}(J)$, while the component in $\ker(J)$ is entirely invisible. Figure 8 illustrates this geometry: the observable output space sees only the $(v_{1},v_{2})$ projection, while $v_{3}\in\ker(J)$ leaves no trace in any output. In language models, $\ker(J_{\ell})$ occupies 86–94% of the ambient activation space, so almost any perturbation to a steering vector goes undetected by behavioral evaluation.
@@
Figure 8: Geometric intuition for null-space ambiguity. The output depends only on the $(v_{1},v_{2})$ projection (row space of $J$), while $v_{3}\in\ker(J)$ is invisible to all outputs.
F.2 Why more data does not resolve the ambiguity
A natural conjecture is that collecting diverse prompts could reduce the null space by probing different output dimensions. Each observation provides $o_{i}=J_{\ell}(x_{i})v+\eta_{i}$, and stacking $N$ observations gives a stacked Jacobian:
$J_{\text{stack}}=\begin{bmatrix}J_{\ell}(x_{1})\\ \vdots\\ J_{\ell}(x_{N})\end{bmatrix}\in\mathbb{R}^{(NV)\times d}$ (36)
with null space $\ker(J_{\text{stack}})=\bigcap_{i=1}^{N}\ker(J_{\ell}(x_{i}))$. In principle, diverse prompts that probe different output dimensions could reduce this intersection. However, directions $v_{0}$ satisfying $J_{\ell}(x)v_{0}=0$ for all $x$ are determined by model weights, not the prompt distribution. When all prompts probe the same effective output subspace, their Jacobians share a common null space and the intersection does not shrink regardless of how many prompts are added. Figure 9 illustrates this: each prompt measures the same active dimensions, leaving the null-space intersection unchanged. The Fisher information along these null-space directions is identically zero, so by the Cramér-Rao bound the estimation variance is unbounded. Non-identifiability is not a small-sample artifact that more data can resolve — it is an intrinsic property of the model geometry.
@@ F.2 Why more data does not resolve the ambiguity
Figure 9: Why prompt diversity does not resolve null-space ambiguity. Each prompt probes the same effective subspace, leaving the null-space intersection unchanged. Numbers reflect Qwen2.5-3B ($d=2048$, rank 139).
F.3 Why orthogonal perturbations preserve semantic coherence
Any perturbation $v_{\perp}\perp v$ decomposes as $v_{\perp}=v_{\perp,\text{row}}+v_{\perp,\text{null}}$, where $v_{\perp,\text{row}}\in\operatorname{row}(J_{\ell})$ is observable through outputs and $v_{\perp,\text{null}}\in\ker(J_{\ell})$ is invisible. For a random $v_{\perp}$, the expected fraction landing in the null space is approximately $\dim(\ker(J_{\ell}))/d\approx 0.86$$0.94$, meaning the bulk of any random perturbation is automatically undetectable. The remaining row-space component contributes an output change $J_{\ell}v_{\perp,\text{row}}$, but this is a random direction in the row space rather than a structured semantic shift. The output change from the original steering vector, $J_{\ell}v$, produces aligned and directional changes across the vocabulary that consistently shift the target trait. The row-space component of the perturbation, by contrast, produces diffuse and incoherent changes that average out across tokens without pushing any semantic dimension in a consistent direction.
To make the distinction concrete: if $J_{\ell}v$ is a strong wind blowing consistently northward, then $J_{\ell}v_{\perp,\text{row}}$ is turbulence. It may carry comparable energy but lacks directional coherence so the overall trajectory remains northward. This is why orthogonal perturbations leave semantic classifier scores nearly unchanged even when their norm is comparable to the steering vector itself.
Appendix G Prompt Construction Details
We construct contrastive prompt pairs for each semantic trait using template-based generation. Table 6 provides concrete examples for all five traits used in the empirical validation. Held-out evaluation prompts are drawn from a separate set of 50 prompts not used in vector extraction, covering in-distribution, topic-shift, and genre-shift conditions for the multi-environment experiment.
Table 6: Example contrastive prompt pairs per trait.
@@ Appendix G Prompt Construction Details
Sentiment Write an enthusiastic review of a product you loved. Write a negative review of a product you disliked.
Truthfulness Provide an accurate factual summary of this event. Make up plausible-sounding details about this event.
Agreeableness Respond warmly and supportively to a colleague’s proposal. Push back sharply on a colleague’s proposal.